OpenStudy (anonymous):

2. The width, w, of a rectangular playground is x+3 . The area of the playground is x^3-7x+6. What is an expression for the length of the playground? I know the answer I just need help working it out?

OpenStudy (anonymous):

(x+3)L = x^3-7x+6

OpenStudy (anonymous):

L is the length

OpenStudy (anonymous):

thank u!!!!!

OpenStudy (anonymous):

factor x^3-7x+6 to get (x-1)(x-2)(x+3)

OpenStudy (anonymous):

(x+3) will cancel

OpenStudy (anonymous):

giving you L = (x-1)(x-2)